OpenStudy (anonymous):

The equation, y = 10(1.0605)^x, represents the winning prize money, in thousands, for first place in the Indianapolis 500 auto race. x represents the number of years after 1911, the first year. What is the value for 1911 (x = 0)? This is the prize money, in thousands. For instance, if you got 23, this represents $23,000. A. $10,605 B. $10,000 C. $0

jimthompson5910 (jim_thompson5910):

\[\Large y = 10(1.0605)^x\] \[\Large y = 10(1.0605)^{0}\] \[\Large y = 10(1)\] \[\Large y = 10\] So the answer is choice B) $10,000